OpenStudy (aravindg):

can anyone answer the question on derivative? can anyone answer the question on derivative? @Mathematics

OpenStudy (anonymous):

whats the question?

OpenStudy (anonymous):

i will try ...

OpenStudy (aravindg):

u will?

OpenStudy (aravindg):

if y= [2/(sqrt a^2-b^2)] x tan inverse [sqrt (a-b/a+b)]tan 1/2 x] then show that dy/dx=1/(a+bcosx)

OpenStudy (aravindg):

can u do it its urgent

OpenStudy (anonymous):

its out of my league, try using globalchautari.com /////////this must help you somehow...

OpenStudy (aravindg):

?wat is that for?

OpenStudy (anonymous):

visit, globalchautari.com, there you will find box for entering mathematical equations........try using it....

OpenStudy (anonymous):

to saruz: what is difference b/w http://www.wolframalpha.com and globalchautari.com?

OpenStudy (perl):

y= [2/(sqrt a^2-b^2)] x tan inverse [sqrt (a-b/a+b)]tan 1/2 x]

OpenStudy (anonymous):

you need to put more parenthesis: is it (a-b)/(a+b) or a-b/a+b as it is?

OpenStudy (anonymous):

So it looks like \[y=\frac{2}{\sqrt{a^2-b^2}} x tan^{-1}\left(\sqrt{\frac{a-b}{a+b}tan\frac{x}{2}}\right)\]?

OpenStudy (aravindg):

no tan 1/2x

OpenStudy (aravindg):

otside square root

OpenStudy (aravindg):

pls help urgent

OpenStudy (anonymous):

\[y=\frac{2}{\sqrt{a^2-b^2}} x tan^{-1}\left(\sqrt{\frac{a-b}{a+b}}\right)tan\frac{1}{2x}\]?

OpenStudy (aravindg):

x in top

OpenStudy (aravindg):

and i doubt if middle is x or into

OpenStudy (aravindg):

helpppppp

OpenStudy (anonymous):

\[y=\frac{2}{\sqrt{a^2-b^2}} x tan^{-1}\left(\sqrt{\frac{a-b}{a+b}}\right)tan(\frac{1}{2}x)\]?

OpenStudy (aravindg):

k

OpenStudy (aravindg):

no

OpenStudy (aravindg):

tan 1/2 x inside bracket of tan inverse

OpenStudy (anonymous):

That is why you need to type it by yourself with enough parenthesis

OpenStudy (anonymous):

\[y=\frac{2}{\sqrt{a^2-b^2}} x tan^{-1}\left(\sqrt{\frac{a-b}{a+b}}tan(\frac{1}{2}x)\right)\]?

OpenStudy (aravindg):

ya

OpenStudy (aravindg):

i thinkk middle is into *

OpenStudy (aravindg):

not x

OpenStudy (anonymous):

What do you mean by 'into'? Division?

OpenStudy (aravindg):

multiply

OpenStudy (aravindg):

*tan inversse

OpenStudy (aravindg):

understand

OpenStudy (anonymous):

Right, once more for absolute clarity then \[y=\frac{2}{\sqrt{a^2-b^2}} \times tan^{-1}\left(\sqrt{\frac{a-b}{a+b}}tan(\frac{1}{2}x)\right)\]

OpenStudy (aravindg):

y a finalized plz giv answr

OpenStudy (anonymous):

@:MarinaDL, there is no any differences with mathematical section , however, there are lots of differences that you have noticed yourself.....

OpenStudy (anonymous):

to saruz: thanks for the answer. It is easier to ask for short summary from somebody experienced than to figure out it on my own.

OpenStudy (anonymous):

Alright, we've got an arctan, a tan, and a multiplication. \[\frac{d}{dx}tan^{-1}(x) = \frac{1}{1+x^2}\]\[\frac{d}{dx}tan(x) = sec^2 x\]and\[\frac{d}{dx} ax = a\] The first fraction is a constant (let's call it k for now) since it doesn't contain an x. The fraction inside the arctan is also a constant, let's call it m. \[k = \frac{2}{\sqrt{a^2 - b^2}}\] \[m=\frac{a-b}{a+b}\] Now we can chain the remaining terms: \[y = k \times tan^{-1}(u) \Rightarrow \frac{dy}{du} = \frac{k}{1+u^2}\]\[u = m \times tan(v) \Rightarrow \frac{du}{dv} = m\times sec^2v\]and\[v=\frac{1}{2}x \Rightarrow \frac{dv}{dx} = \frac{1}{2}\] \[\frac{dy}{dx}=\frac{dy}{du}\times\frac{du}{dv}\times\frac{dv}{dx}\] Is all of that cool AravindG?

OpenStudy (anonymous):

\[\frac{dy}{dx}=\frac{dy}{du}\times\frac{du}{dv}\times\frac{dv}{dx}\]

OpenStudy (anonymous):

The question is to 'show that \(\frac{dy}{dx}=\frac{1}{a+bcosx} \)', so that is hopefully the answer.

OpenStudy (anonymous):

You should probably work through the actual algebra at the end, and make sure you understand why I did all that chaining.

OpenStudy (aravindg):

x=sin^3t/sqrt(cos 2t),y=cos^3t/sqrt(cos 2t) at t= pi/6 find dy/dx

OpenStudy (anonymous):

y does not seem to be a function of x, so dy/dx would be 0.

OpenStudy (anonymous):

I guess you could do some substitution, \[x=\frac{sin^3t}{\sqrt{cos(2t)}}\Rightarrow \sqrt{cos(2t)}=\frac{sin^3t}{x}\]\[y=\frac{cos^3t}{\sqrt{cos(2t)}} \Rightarrow y=\frac{cos^3t}{\frac{sin^3t}{x}}=\frac{x cos^3t}{sin^3t}=\frac{x}{tan^3t}\]I think that the tan is constant with respect to x, so \[\frac{dy}{dx} = \frac{1}{tan^3t}\]

OpenStudy (aravindg):

wow now find values of a.,b,c such that lim x tends to 0 ((x(a+bcosx)-c*sinx)/x^5)=1

OpenStudy (perl):

you have a funny expression inside the arctan

OpenStudy (perl):

arctan ( k * tan u )

OpenStudy (anonymous):

I dunno much about limits I'm afraid!

OpenStudy (aravindg):

wt abt this find derivativre tan inverse (5 a x/a^2-6 x^2)

OpenStudy (anonymous):

You're going to have to use the chain rule, and the quotient rule on that one.

OpenStudy (aravindg):

plz show u see its very very urgent

OpenStudy (perl):

ok

OpenStudy (perl):

let k = sqrt ( a-b) / sqrt ( a+ b )

OpenStudy (perl):

y = 2k/ (a-b) arctan [ k tan (1/2x)]

OpenStudy (perl):

y' = 2k/(a-b) * 1/ { 1 + ( ktan(1/2*x))^2}* sec^2 ( 1/2*x)*1/2

OpenStudy (anonymous):

\[y=tan^{-1}\left(\frac{5ax}{a^2-6x^2}\right)\] Rewrite it as \[y=tan^{-1}(w) \Rightarrow \frac{dy}{dx} = \frac{1}{1+w^2}\]\[w=\frac{5ax}{a^2-6x^2}\Rightarrow \frac{dy}{dx}=\frac{v \times \frac{du}{dx} - u\times \frac{dv}{dx}}{v^2}\] Where \(u = 5ax, v = a^2 - 6x^2\) I'm sure you can handle it from there bro!

OpenStudy (anonymous):

Sorry, that first dy/dx should be dy/dw.

OpenStudy (anonymous):

And the second one is dw/dx.

OpenStudy (perl):

y' = 2k/(a-b) * 1/ { 1 + ( ktan(1/2*x))^2}* sec^2 ( 1/2*x)*1/2 y' = 2k/(a-b) * 1/ (1 + k^2*tan^2(1/2*x)* 1/cos^2( 1/2*x)*1/2

OpenStudy (aravindg):

find dy/dx y=(x+1)^2sqrt(x-1)/(x+4)^3 e^x

OpenStudy (perl):

are you doing a different problem now ?

OpenStudy (aravindg):

((x+1)^2sqrt(x-1))/((x+4)^3 e^x)

OpenStudy (aravindg):

u cntinue i need al answers

OpenStudy (perl):

y = 2k/ (a-b) arctan [ k tan (1/2x)] y' = 2k/(a-b) * 1/ { 1 + ( ktan(1/2*x))^2}* k*sec^2 ( 1/2*x)*1/2

OpenStudy (perl):

y' = 2k/(a-b) * 1/ (1 + k^2*tan^2(1/2*x)* k/cos^2( 1/2*x)*1/2 y' = k/(a-b) * 1/ [k*cos^2(1/2x) + k^3*sin^2(1/2*x)]

OpenStudy (perl):

y' = 1/(a-b) * 1/ [ cos^2(1/2x) + k^2 sin^2 (1/2*x)

OpenStudy (perl):

y' = 1/(a-b) * 1/ ( cos^2(1/2x) + k^2 [1 - cos^2 (1/2*x)] )

OpenStudy (perl):

still there?

OpenStudy (perl):

ok i got it , it took a while, wow